# A Survey on Failure Analysis and Fault Injection in AI Systems

### v1

@@ Abstract.
The rapid advancement of Artificial Intelligence (AI) has led to its integration into various areas, especially with Large Language Models (LLMs) significantly enhancing capabilities in Artificial Intelligence Generated Content (AIGC). However, the complexity of AI systems has also exposed their vulnerabilities, necessitating robust methods for failure analysis (FA) and fault injection (FI) to ensure resilience and reliability. Despite the importance of these techniques, there lacks a comprehensive review of FA and FI methodologies in AI systems. This study fills this gap by presenting a detailed survey of existing FA and FI approaches across six layers of AI systems. We systematically analyze 160 papers and repositories to answer three research questions including (1) what are the prevalent failures in AI systems, (2) what types of faults can current FI tools simulate, (3) what gaps exist between the simulated faults and real-world failures. Our findings reveal a taxonomy of AI system failures, assess the capabilities of existing FI tools, and highlight discrepancies between real-world and simulated failures. Moreover, this survey contributes to the field by providing a framework for fault diagnosis, evaluating the state-of-the-art in FI, and identifying areas for improvement in FI techniques to enhance the resilience of AI systems.
Failure Analysis, Fault Injection, Chaos Engineering, AI system, MLOps
copyright: acmcopyrightjournalyear: 2024journal: CSURdoi: XXXXXXX.XXXXXXXcopyright: noneccs: General and reference Surveys and overviewsccs: General and reference Reliabilityccs: General and reference Performanceccs: Software and its engineering Software organization and propertiesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Artificial Intelligence (AI) has made significant strides over the past decade, permeating both academic and industrial areas. Large Language Models (LLMs), in particular, have proven to be a game changer, propelling AI to unprecedented heights and facilitating a myriad of applications in fields such as software engineering (Pearce et al., 2023; Wei et al., 2023; Xia et al., 2023) and human language translation (Kocmi and Federmann, 2023; Hendy et al., 2023; Bang, 2023). This evolution has led to the integration of AI models into a growing array of products, transforming them into sophisticated AI systems. Notable examples of such integration include Gemini (Google, 2024), Bing (Microsoft, 2024), and ChatGPT (OpenAI, 2024a), which underscore the pivotal role of AI in enhancing and expanding the capabilities of modern technology solutions.
@@
The escalating complexity and ubiquity of AI systems necessitate addressing their inherent vulnerabilities and failure-related challenges. A Meta AI report (Zhang et al., 2022) points out over 100 failures during the training of OPT-175B. Similarly, ChatGPT encountered 173 outages in 2023, causing a maximum user impact over 427 minutes (OpenAI, 2024b). Such failures can degrade user experience, and even incur financial losses. Hence, mitigating AI system failures is of paramount importance.
@@
Failure analysis (FA) and fault injection (FI) techniques are instrumental in identifying limitations and bolstering the reliability of AI systems. Researchers and practitioners alike have embarked on extensive investigations into AI system failures. Studies (Liu et al., 2023b; Chen et al., 2023; Zhang et al., 2018; Makkouk et al., 2022; Tambon et al., 2024; Li et al., 2023; Islam et al., 2019; Humbatova et al., 2020) have analyzed AI system failures from platforms like Stack Overflow or Github, while others (Jeon et al., 2019; Zhang et al., 2020b; Gao et al., 2023b; Ostrouchov et al., 2020; Gupta et al., 2017; Taherin et al., 2021; Tiwari et al., 2015a) have focused on failures in large-scale production AI systems. Such failure analyses enable the identification of patterns, root causes, and locations, thereby informing FI techniques. FI, a proactive approach, uncovers system weaknesses on resiliency before they become catastrophic failures. By deliberately injecting faults or abnormal conditions into systems, teams can evaluate and enhance their resilience to unexpected disruptions. Some existing FI approaches (Chan et al., 2022, 2021; Ma et al., 2018; Sambasivan et al., 2021; Hendrycks and Dietterich, 2019; Jia et al., 2022; Humbatova et al., 2021; Shen et al., 2018) mimic faults in AI systems engineered by humans, while others (Zhang et al., 2019b; Li et al., 2017; Oliveira et al., 2017; Salami et al., 2018; Zhang et al., 2019a; Deng et al., 2015a) simulate hardware errors.
@@
Despite the progress, a comprehensive survey on FA and FI in AI systems is conspicuously absent. Furthermore, a gap exists between FI and FA, resulting in insufficient consideration of FA outcomes when crafting FI tools. Therefore, this study presents a comprehensive survey aimed at exploring and evaluating existing research for FA and FI in AI systems. We have meticulously reviewed and analyzed 160 corresponding papers and code repositories. As shown in Fig. 1, an AI system typically comprises six layers - AI Service, AI Model, AI Framework, AI Toolkit, AI Platform, and AI Infrastructure (Tsaih et al., 2023). We attempt to address three research questions at each layer as follows.
@@
  • RQ2: What types of faults can current FI tools simulate?
  • RQ3: What gaps do exist between the simulated faults and the real-world failures?
RQ1 aims to catalog and analyze the failures that have occurred in current AI systems. Understanding these failures is crucial for several reasons. Since it helps in identifying common vulnerabilities within AI systems, informs developers about potential areas of improvement, and contributes to the development of more reliable AI applications. RQ2 explores the capabilities of existing FI tools designed for AI systems. The ability to simulate a wide range of faults is essential for evaluating and enhancing the robustness and fault tolerance of AI systems. RQ3 investigates the gap between simulated faults and real-world AI system failures, aiming to understand the limitations of current FI tools in producing the full spectrum of potential failures. Moreover, understanding these gaps helps in improving FI tools, and ultimately contributes to develop more resilient AI systems.
By examining the current landscape and identifying critical research gaps, this survey provides valuable insights for researchers and practitioners working towards building reliable and resilient AI systems. This study makes the following contributions:
  • We present a comprehensive analysis and taxonomy of failures occurring at different layers of AI systems. By systematically characterizing these failures, we provide a valuable framework that can serve as a reference for failure diagnosis in AI systems.
  • We conduct an in-depth examination of the capabilities of existing FI tools across various layers of AI systems. We offer insights into the state-of-the-art in simulating and reproducing potential failures. This work provides a foundation for assessing the reliability of AI systems.
  • We explore the discrepancies between FI tools and real-world AI system failures. We identify the limitations of current FI approaches in simulating potential failure scenarios. By shedding light on these gaps, we emphasize the need for more comprehensive FI techniques in AI systems.
The rest of this paper is organized as follows. Section 2 provides background information on FA and FI in AI systems, followed by Section 3, which outlines our systematic literature review methodology. The subsequent sections analyze FA and FI in different layers of AI systems, including the AI service layer (Section 4), AI model layer (Section 5), AI framework layer (Section 6), AI toolkit layer (Section 7), AI platform layer (Section 8), and AI infrastructure layer (Section 9). Section 10 highlights research opportunities about FI in AI systems. The article concludes in Section 11.
@@ 2.1. Failures and Faults
We adopt the definitions of failures and faults proposed by previous work (Avizienis et al., 2004; Salfner et al., 2010). Furthermore, we provide additional extensions and interpretations specific to AI systems.
  • Failure is defined as ”an incident that occurs when the delivered service deviates from the correct service” (Avizienis et al., 2004). In the context of AI systems, failures can manifest in various ways. For example, a failure can occur when AI services become unreachable, and when the behavior of AI services does not meet the expected outcome (e.g., generating semantically incorrect text). These failures indicate a deviation from the desired or expected behavior of the AI system.
@@ 2.2. Failure Comparisons between AI and Cloud System
As illustrated in Figure 2, while both AI and cloud systems are susceptible to common failures such as power disruptions and network outages, certain faults are unique or more critically impactful to AI systems. For instance, GPU failures, which might be relatively inconsequential in traditional cloud environments, can severely affect the performance and availability of AI systems. This distinction underscores the importance of conducting a comprehensive survey on FA and FI specifically for AI systems, especially in the current era dominated by ”Large Models”. This need forms one of the primary motivations behind the research presented in this article.
@@ 2.3. Failure Analysis and Fault Injection
FI is a widely adopted technique for assessing and improving the reliability and security of systems, including AI systems. It involves the deliberate introduction of faults into a system to observe its behavior and validate its fault tolerance mechanisms. FI can be applied in various forms, such as software fault injection (e.g., mutation test (Jia and Harman, 2011) and API interception (Yu et al., 2023a; Wan et al., 2021)) or hardware fault injection (e.g., simulating hardware failures (Nie et al., 2018) and environmental disturbances (Khan et al., 2023)).
@@
Consequently, FA and FI form closely intertwined processes that significantly contribute to the assessment and enhancement of AI system reliability. The insights derived from FA guide the selection and design of FI scenarios. The iterative feedback loop established by fault analysis and FI facilitates the continuous improvement of AI systems, thereby serving as the other motivation driving the research presented in this article.
@@ 3. Survey Methodology
To systematically collect the publications for conducting this survey, we specially constructed and maintained a repository  [^1] https://github.com/IntelligentDDS/awesome-papers/tree/main/Fault_tolerance#ai-system about FA and FI in AI systems. We first searched relevant papers in online digital libraries and extended the repository by manually inspecting all references of these papers.
To begin with, we searched several popular online digital libraries (e.g., IEEE Xplore, ACM Digital Library, Springer, Elsevier, and Wiley) with the following keywords: ”failure and machine learning”, ”failure and deep learning”, ”failure and AI”, ”fault injection and machine learning”, ”fault injection and deep learning”, ”fault injection and AI”, etc. We mainly focused on regular papers published relevant conferences (e.g., ICSE, FSE, ASE, SOSP, OSDI, ATC, NSDI, SC, DSN, ISSRE, AAAI, ICML, NeurIPS, etc) and journals (e.g., TSE, TOSEM, EMSE, TDSC, TPDS, TSC, etc). Upon identifying relevant papers, we conducted a recursive examination of the references, which allowed us to manually inspect each reference and code repository of these papers. This process enabled us to collect additional publications and repositories related to our survey topics.

### v2

@@ Abstract.
The rapid advancement of Artificial Intelligence (AI) has led to its integration into various areas, especially with Large Language Models (LLMs) significantly enhancing capabilities in Artificial Intelligence Generated Content (AIGC). However, the complexity of AI systems has also exposed their vulnerabilities, necessitating robust methods for failure analysis (FA) and fault injection (FI) to ensure resilience and reliability. Despite the importance of these techniques, there lacks a comprehensive review of FA and FI methodologies in AI systems. This study fills this gap by presenting a detailed survey of existing FA and FI approaches across six layers of AI systems. We systematically analyze 142 studies to answer three research questions including (1) what are the prevalent failures in AI systems, (2) what types of faults can current FI tools simulate, (3) what gaps exist between the simulated faults and real-world failures. Our findings reveal a taxonomy of AI system failures, assess the capabilities of existing FI tools, and highlight discrepancies between real-world and simulated failures. Moreover, this survey contributes to the field by providing a framework for fault diagnosis, evaluating the state-of-the-art in FI, and identifying areas for improvement in FI techniques to enhance the resilience of AI systems.
Failure Analysis, Fault Injection, Chaos Engineering, AI system, MLOps, LLMOps
copyright: acmcopyrightjournalyear: 2025journal: TOSEMdoi: XXXXXXX.XXXXXXXccs: General and reference Surveys and overviewsccs: General and reference Reliabilityccs: General and reference Performanceccs: Software and its engineering Software organization and propertiesccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Artificial intelligence (AI) has made significant strides in the past decade, permeating both academic and industrial areas. Large Language Models (LLMs), in particular, have proven to be a game changer, propelling AI to unprecedented heights and facilitating a myriad of applications in fields such as software engineering (Pearce et al., 2023; Wei et al., 2023; Xia et al., 2023) and human language translation (Kocmi and Federmann, 2023; Hendy et al., 2023; Bang, 2023). This evolution has led to the integration of AI models into a growing array of products, which transform them into sophisticated AI systems. Notable examples of such integration include Gemini (Google, 2024), Bing (Microsoft, 2024), and ChatGPT (OpenAI, 2024a), which underscore the pivotal role of AI in enhancing and expanding the capabilities of modern technology solutions.
@@
The increasing complexity and ubiquity of AI systems necessitate addressing their inherent vulnerabilities and failure-related challenges. A Meta AI report (Zhang et al., 2022) points out more than 100 failures during OPT-175B training. Similarly, ChatGPT encountered 173 outages in 2023, causing a maximum user impact of 427 minutes (OpenAI, 2024b). Such failures can degrade user experience and even incur financial losses. Hence, mitigating AI system failures is of paramount importance.
The real-world implications of failure analysis in AI systems are far-reaching. They not only affect the performance and reliability of AI systems, but also have a profound impact on the industries and sectors that rely on these systems. For example, in healthcare, a failure in an AI system could lead to incorrect diagnoses or treatment recommendations, potentially endangering lives. In the financial sector, AI system failures could result in significant financial losses due to erroneous transactions or predictions. Therefore, understanding and addressing these failures is not just a technical challenge but a social imperative.
@@
Failure analysis (FA) and fault injection (FI) techniques are instrumental in identifying limitations and bolstering the reliability of AI systems. Researchers and practitioners alike have conducted extensive investigations into AI system failures. Studies (Liu et al., 2023b; Chen et al., 2023; Zhang et al., 2018; Makkouk et al., 2022; Tambon et al., 2024; Li et al., 2023; Islam et al., 2019; Humbatova et al., 2020) have analyzed AI system failures from platforms such as Stack Overflow or Github, while others (Jeon et al., 2019; Zhang et al., 2020b; Gao et al., 2023b; Ostrouchov et al., 2020; Gupta et al., 2017; Taherin et al., 2021; Tiwari et al., 2015a) have focused on failures in large-scale production AI systems. Such failure analyses enable the identification of patterns, root causes, and locations, thereby informing FI techniques. FI, a proactive approach, uncovers system weaknesses in resilience before they become catastrophic failures. By deliberately injecting faults or abnormal conditions into systems, teams can evaluate and enhance their resilience to unexpected disruptions. Some existing FI approaches (Chan et al., 2022, 2021; Ma et al., 2018; Sambasivan et al., 2021; Hendrycks and Dietterich, 2019; Jia et al., 2022; Humbatova et al., 2021; Shen et al., 2018) mimic faults in AI systems engineered by humans, while others (Zhang et al., 2019b; Li et al., 2017; Oliveira et al., 2017; Salami et al., 2018; Zhang et al., 2019a; Deng et al., 2015) simulate hardware errors.
@@
Despite progress, a comprehensive survey on FA and FI in AI systems is conspicuously absent. Furthermore, there is a gap between FI and FA, resulting in insufficient consideration of FA outcomes when designing FI tools. Therefore, this study presents a comprehensive survey aimed at exploring and evaluating existing research on FA and FI in AI systems. We meticulously reviewed and analyzed 142 corresponding publications. Motivated by previous studies (Tsaih et al., 2023), we divide an AI system into six layers including AI Service, AI Model, AI Framework, AI Toolkit, AI Platform, and AI Infrastructure. The relationships between these layers are illustrated in Fig. 1, while detailed explanations are provided in Table 1. We attempt to address three research questions at each layer as follows.
@@
Table 1. Definition of layers in AI system.
Stack Definition Representative
AI Service Well-defined AI-enabled services that can be fulfilled by API function call request. ChatGPT
AI Model AI-related models that power the AI capabilities of the system GPT-4
AI Framework AI-related frameworks for defining AI/ML architectures and invoking algorithms, library, and accelerator drive on the hardware. Pytorch
AI Tookit An intermediate interface between the AI framework and underlying devices CUDA
AI Platform AI-related platform that enable large-scale organizations to design, develop, deploy and operate AI services Ray
AI Infrastructure Hysical and virtualized resources necessary for the deployment and operation of AI services. GPU, TPU
@@
  • RQ2: What types of fault can current FI tools simulate?
  • RQ3: What gaps exist between simulated faults and real-world failures?
RQ1 aims to catalog and analyze the failures that have occurred in current AI systems. Understanding these failures is crucial for several reasons. Since it helps to identify common vulnerabilities within AI systems, it informs developers about potential areas of improvement and contributes to the development of more reliable AI applications. RQ2 explores the capabilities of existing FI tools designed for AI systems. The ability to simulate a wide range of faults is essential for evaluating and improving the robustness and fault tolerance of AI systems. RQ3 investigates the gap between simulated faults and real-world AI system failures, with the aim of understanding the limitations of current FI tools in producing the full spectrum of potential failures. Moreover, understanding these gaps helps to improve FI tools and ultimately contributes to develop more resilient AI systems.
By examining the current landscape and identifying critical research gaps, this survey provides valuable insights for researchers and practitioners working toward building reliable and resilient AI systems. This study makes the following contributions.
  • We present a comprehensive analysis and taxonomy of failures that occur in different layers of AI systems. By systematically characterizing these failures, we provide a valuable framework that can serve as a reference for failure diagnosis in AI systems. This taxonomy paves the way for future research by highlighting specific areas where AI systems are most vulnerable, thereby guiding targeted improvements and innovations.
  • We conduct an in-depth examination of the capabilities of existing FI tools across various layers of AI systems. We offer insights into the state-of-the-art in simulating and reproducing potential failures. This work provides a foundation for assessing the reliability of AI systems. Our findings underscore the importance of developing more sophisticated FI tools that can better mimic real-world failure scenarios, thus enhancing the robustness of AI systems.
  • We explore the discrepancies between FI tools and real-world AI system failures. We identify the limitations of current FI approaches in simulating potential failure scenarios. By shedding light on these gaps, we emphasize the need for more comprehensive FI techniques in AI systems. Addressing these gaps is crucial for advancing the field, as it will lead to the creation of more resilient AI systems capable of withstanding a broader range of failure conditions.
The remainder of this paper is organized as follows. Section 2 provides background information on FA and FI in AI systems, followed by Section 3, which outlines our systematic literature review methodology. The subsequent sections analyze FA and FI in different layers of AI systems, including the AI service layer (Section 4), AI model layer (Section 5), AI framework layer (Section 6), AI toolkit layer (Section 7), AI platform layer (Section 8), and AI infrastructure layer (Section 9). Section 10 outlines potential threats to validity in our study. Section 11 highlights research opportunities on FI in AI systems. The article concludes in Section 12.
@@ 2.1. Failures and Faults
We adopt the definitions of failures and faults proposed in previous work (Avizienis et al., 2004; Salfner et al., 2010). Furthermore, we provide additional extensions and interpretations specific to AI systems.
  • Failure is defined as “an incident that occurs when the delivered service deviates from the correct service”  (Avizienis et al., 2004). In the context of AI systems, failures can manifest in various ways. For example, a failure can occur when AI services become unreachable and when the behavior of AI services does not meet the expected outcome (e.g., generating semantically incorrect text). These failures indicate a deviation from the desired or expected behavior of the AI system.
@@ 2.2. Failure Comparisons between AI and Cloud System
As illustrated in Fig. 2, while AI and cloud systems are susceptible to common failures such as power disruptions and network outages, certain faults are unique or more critical to AI systems. For instance, GPU failures, which might be relatively inconsequential in traditional cloud environments, can severely affect the performance and availability of AI systems. This distinction underscores the importance of conducting a comprehensive survey on FA and FI specifically for AI systems, especially in the current era dominated by “Large Models”. This need forms one of the primary motivations behind the research presented in this article.
@@ 2.3. Failure Analysis and Fault Injection
Figure 3. Overall life cycle of an AI system failure.
@@
FI is a widely adopted technique to assess and improve the reliability and security of systems, including AI systems. It involves the deliberate introduction of faults into a system to observe its behavior and validate its fault tolerance mechanisms. FI can be applied in various forms, such as software fault injection (e.g., mutation test (Jia and Harman, 2011) and API interception (Yu et al., 2023a; Wan et al., 2021)) or hardware fault injection (e.g., simulating hardware failures (Nie et al., 2018) and environmental disturbances (Khan et al., 2023)).
@@
Consequently, FA and FI form closely intertwined processes that contribute significantly to the evaluation and enhancement of AI system reliability. The insights derived from FA guide the selection and design of FI scenarios. The iterative feedback loop established by fault analysis and FI facilitates the continuous improvement of AI systems, thereby serving as the other motivation driving the research presented in this article.
@@ 3. Survey Methodology
Figure 4. Search strategy of our study.
3.1. Search Strategy
We visualize our research process in Fig. 4 and describe the different steps in this section. Our study employs a hybrid search strategy (Mourão et al., 2020), combining a systematic search in Scopus  [^1] https://www.scopus.com/ with a bidirectional snowballing approach (referred to as “Scopus + BS*FS”) (Wohlin, 2014). This strategy first executes a search query in Scopus to obtain an initial set of papers, which serves as the seed set for subsequent snowballing. Scopus was chosen as the primary database due to its extensive coverage of peer-reviewed publications from top software engineering journals and conferences, including IEEE Xplore, ACM Digital Library, ScienceDirect (Elsevier), and Springer. Previous studies (Costal et al., 2021) have demonstrated that Scopus provides optimal coverage compared to other databases.
The search string was iteratively refined through discussions among the authors, ensuring it captured relevant studies while minimizing noise. The final search string executed in September 2024 was as follows:
( TITLE (Failure OR Fault OR "Fault Injection" OR "Bug" OR "Error" )
AND
TITLE-ABS-KEY ( "AI System" OR "AI Application"
OR "Artificial Intelligence System" OR "Artificial Intelligence Application"
OR "ML System" OR "ML Software"
OR "Machine Learning System" OR "Machine Learning Application"
OR "Deep Learning System" OR "Deep Learning Application"
OR "DL System" OR "DL Application"
OR "Large Language Model System" OR "Large Language Model Application"
OR "LLM System" OR "LLM Application")
AND
PUBYEAR > 1999 AND PUBYEAR < 2025
AND
(LIMIT-TO ( SUBJAREA , "COMP" ) )
This query yielded 331 potentially relevant studies. Each study was independently screened by two researchers using predefined inclusion and exclusion criteria (detailed in Section 3.1.1). After screening, 38 studies formed the seed set for snowballing. We conducted two rounds of bidirectional snowballing (Section 3.1.2), resulting in a final set of 142 publications on failure analysis (FA) and fault injection (FI) in AI systems, spanning the years 2001 to 2024.
3.1.1. Screening of Studies
The following inclusion (I) and exclusion (E) criteria were applied during the selection of primary studies:
  • Inclusion criteria
    • IC1: The study primarily focuses on failure analysis for AI systems.
    • IC2: The study primarily focuses on fault injection for AI systems.
  • Exclusion criteria
    • EC1: The study does not fulfill IC1 or IC2.
    • EC2: The study is not accessible, even after contacting the authors.
    • EC3: The study is not entirely written in English.
    • EC4: The study is a pure secondary or tertiary study.
    • EC5: The study was published before January 2000 or after September 2024.
A study had to meet all inclusion criteria and none of the exclusion criteria to be included in the primary study set. To ensure consistency, a pilot selection process was conducted, where three researchers with expertise in AI systems and software engineering independently evaluated five studies and discussed their decisions in a consensus meeting. After refining the selection process, it was applied to all potentially relevant studies. Each study was independently evaluated by two researchers, and consensus was required for final inclusion. This process resulted in 38 studies forming the seed set for snowballing.
3.1.2. Snowballing Approach
To comprehensively identify relevant studies, we applied both backward and forward snowballing, as recommended by Wohlin et al. (Wohlin, 2014). In this process, all papers that either cited or were cited by a paper in the seed set were evaluated for inclusion in the final set of primary studies. The same inclusion and exclusion criteria and screening process were applied, with each study independently evaluated by two researchers and consensus required for inclusion. After the first round of bidirectional snowballing, 44 additional studies were included. A second round of bidirectional snowballing added 60 more studies, resulting in a final set of 142 publications.
3.2. Data Extraction and Mapping Process
To address the research questions (RQ1–RQ3) and systematically analyze the relationships between failure analysis (FA) and fault injection (FI) in AI systems, we developed a structured data extraction and mapping process. This process was designed to ensure consistency, reproducibility, and alignment with the six-layer AI system architecture (AI Service, AI Model, AI Framework, AI Toolkit, AI Platform, and AI Infrastructure) illustrated in Fig. 1 and detailed in Table 1.
3.2.1. Data Extraction Framework
We defined a data extraction framework to systematically collect and organize relevant information from the 142 primary studies. The framework included the following key data points for each study:
  • Study Metadata: Publication year, venue, and type (e.g., journal, conference).
  • AI System Layer: The layer(s) of the AI system addressed in the study (e.g., AI Model, AI Service).
  • Failure Analysis (FA): Types of failures reported, failure causes, and their impact on the AI system.
  • Fault Injection (FI): Types of faults simulated, FI techniques used, and their alignment with real-world failures.
  • Gaps Between FA and FI: Identified discrepancies between simulated faults and real-world failures, as well as limitations of current FI tools.
3.2.2. Mapping Process
The extracted data was mapped to the six-layer AI system architecture to address the research questions at each layer. The mapping process involved the following steps:
  1. (1)
    Classification by Layer: Each study was classified into one or more layers based on its primary focus. For example, a study discussing failures in AI model training was classified under the AI Model layer.
  2. (2)
    Iterative Validation: Two authors independently classified the studies, and disagreements were resolved through consensus meetings. This iterative process ensured accuracy and consistency in the classification.
  3. (3)
    Cross-Layer Analysis: Studies addressing multiple layers were analyzed to identify inter-layer dependencies and interactions. For instance, failures in the AI Service layer were often linked to underlying issues in the AI Model or AI Infrastructure layers.
3.2.3. Addressing Research Questions
The extracted and mapped data was analyzed to answer the three research questions:
  • RQ1 (Prevalent Failures): We cataloged and analyzed failures reported in each layer, identifying common vulnerabilities and their root causes. This analysis provides insights into the types of failures that occur in real-world AI systems and highlights areas for improvement.
  • RQ2 (FI Tool Capabilities): We evaluated the types of faults simulated by existing FI tools, assessing their coverage and effectiveness in testing AI system robustness. This analysis reveals the strengths and limitations of current FI tools.
  • RQ3 (Gaps Between FA and FI): We identified discrepancies between simulated faults and real-world failures, uncovering limitations in current FI tools. This analysis informs the design of more effective FI tools that better reflect real-world failure scenarios.
3.2.4. Quality Assurance
To ensure the reliability and validity of the data extraction and mapping process, we implemented the following quality assurance measures:
  • Inter-Rater Reliability: Two authors independently extracted and mapped data for a subset of studies, and the inter-rater agreement was measured using Cohen’s Kappa coefficient (Cohen, 1960). The inter-rater reliability for classification was measured using Cohen’s Kappa (k = 0.82), indicating high consistency. Disagreements were resolved through discussion and consensus.
  • Iterative Refinement: The data extraction framework and mapping process were refined iteratively based on feedback from the authors and preliminary findings. This refinement ensured the framework captured all relevant data points and addressed the research questions effectively.
Through this rigorous data extraction and mapping process, we systematically analyzed the relationships between FA and FI in AI systems, providing a comprehensive understanding of the gaps and opportunities for future research.
3.3. Summary of Our Survey
In total, our recursive search methodology allowed us to collect 142 studies on FA and FI in AI systems, covering 2001 to 2024. Figure 5 shows the histogram of annual studies during the period. We can FA and FI in AI systems has been continuously and actively investigated in the past decades. In particular, we can observe a steady growth in the number of studies since 2015, indicating that the AI systems field has attracted an increasing amount of interest since then.